OpenStudy (anonymous):

Solve for x. (8 2/5 + 7 2/3) + 4 1/3 = x + 13

OpenStudy (anonymous):

A. 6 2/5 B. 7 2/5 C. 8 2/5 D. 16 3/5

OpenStudy (xhawaiiangirlx):

would u like steps or straight answer?

OpenStudy (anonymous):

steps

OpenStudy (xhawaiiangirlx):

okay

OpenStudy (xhawaiiangirlx):

first we need to refine the equations otherwise simplyfiy

OpenStudy (xhawaiiangirlx):

x+13=811/15

OpenStudy (xhawaiiangirlx):

then multiply each side by 15

OpenStudy (xhawaiiangirlx):

15x+15*13=15 811/15

OpenStudy (xhawaiiangirlx):

than refine to 15x+195=811

OpenStudy (xhawaiiangirlx):

then subtract 195 by each side

OpenStudy (xhawaiiangirlx):

15x+195-195=811-195 15x=616

OpenStudy (xhawaiiangirlx):

finally divide both sides by 15 15x/15=616/15

OpenStudy (xhawaiiangirlx):

which the answer 616/15